OpenStudy (anonymous):

integrate tan(^-1)x dx

OpenStudy (goten77):

http://www.wolframalpha.com/input/?i=integrate+tan^-1%28x%29&lk=4 personally i just recognize the chain rule basic setup and apply that \[\int\limits_{?}^{?} \tan ^{-1}u du = -\ln \left| \cos u \right| +c \] the proof takes a while to show but just go by the parent function/chain rule

OpenStudy (anonymous):

thank you